OpenStudy (anonymous):

Use implicit differentiation to find dy/dx given that x^3y-sin2y=4+y^3

OpenStudy (anonymous):

(x^3y)'=3dy/dx x^(3y-1) (-sin2y)'=-2dy/dx cos2y (4+y^3)'=3dy/dx y^2 factor out dy/dx and you will have your answer

OpenStudy (anonymous):

\[x^3y -sin(2y)=4+y^3\] \[3x^2y+x^3y'-2\cos(y)y'=3y^2y'\] \[3x^2y=3y^2y'+2\cos(2y)y'-x^3y'\] \[3x^2y=y'(3y^2+2\cos(2y)-x^3)\] \[\frac{3x^2y}{3y^2+2\cos(2y)-x^3}=y'\]

OpenStudy (anonymous):

check the algebra!

OpenStudy (anonymous):

thats what i have satellite thanks :D